# DreamTalk: When Expressive Talking Head Generation Meets Diffusion Probabilistic Models

### v1

@@ 3.3 Training and Inference
instead of Equation 2. This approach enables controlling the effectiveness of the style reference $\bm{R}$ through adjustment of the scale factor $\omega$.
@@
We utilize PIRenderer [52] as the renderer and meticulously fine-tune it to empower the renderer with emotional expression generation capabilities.
Inference. Our method enables the specification of speaking styles using either reference videos or solely through input audio and portrait. In the case of reference videos, style codes are derived using the style encoder in the denoising network. When relying solely on input audio and portrait, these inputs are processed by the style predictor, which employs a denoising procedure to obtain the style code.
With the style code, the denoising network utilizes the sampling algorithm of DDPM [26] to produce face motions. It first samples a random motion $\bm{m}_{(T)}^{*}\sim{}\mathcal{N}(\bm{0},\bm{I})$ then computes denoised sequences $\{\bm{m}_{(t)}^{*}\},{t=T-1,\dots,0}$ by incrementally removing the noise from $\bm{m}_{(t)}^{*}$. Finally, the motion $\bm{m}_{(0)}^{*}$ is the generated face motion. The sampling process can be accelerated by leveraging DDIM [63]. The output face motions are then rendered into videos by the renderer PIRenderer.
4 Experiments
4.1 Experimental Setup
@@
Baselines. We compare our method with previous methods including: MakeitTalk [105], Wav2Lip [49], PC-AVS [104], AVCT [84], GC-AVT [37], EAMM [30], StyleTalk [46], DiffTalk [58], SadTalker [100], PD-FGC [78], and EAT [20]. For DiffTalk, since the released model is incomplete and unable to generate reasonable results until submission, we perform qualitative comparisons using videos from its released demo. For other methods, we generate the samples using released models or with the help of the authors.
Metrics. We utilize widely used metrics: SSIM [88], the Cumulative Probability of Blur Detection (CPBD) [48], the SyncNet confidence score ($\text{Sync}_{\text{conf}}$[9], the Landmark Distance around mouth area (M-LMD) [6], the Landmark Distance on the full face (F-LMD).
Figure 3: Qualitative comparisons with previous methods.
4.2 Main Results
@@
Quanitative Comparisons. As shown in Tab. 1, our method outperforms previous methods across most metrics. Wav2Lip’s training with SyncNet as a discriminator explains its high SyncNet confidence score, even surpassing the ground truth. Notably, our method’s SyncNet confidence score closely aligns with the ground truth, and it achieves the best M-LMD scores, which indicates its capability for precise lip synchronization. Furthermore, our superior performance in the F-LMD metric demonstrates our method’s proficiency in generating facial expressions consistent with the reference speaking style.
Qualitative comparisons. Fig. 3 shows the qualitative comparisons. The portraits, style references, and audio are all unseen during training. It can be seen that MakeItTalk and AVCT struggle with accurate lip synchronization. While Wav2Lip and PC-AVS synchronize lips accurately, their outputs appear blurry. SadTalker, on the other hand, generally aligns lip movements with audio but occasionally displays unnatural jitters.
@@
EAT’s capability is limited to generating discrete emotions, lacking the finesse for nuanced expressions. For example, in the left case, the style reference shows the speaker narrowing his eyes, but EAT merely produces a generic disgusted look with wide-open glaring eyes. Additionally, as shown in the right case, EAT struggles to maintain a consistent face shape during speaker head movements.
EAMM, GC-AVT, StyleTalk, and PD-FGC demonstrate the ability to produce fine-grained expressions. However, EAMM falls short in lip synchronization, GC-AVT and PD-FGC struggle with preserving speaker identity, and all three have issues rendering a plausible background. We observed that StyleTalk, while capable of generating nuanced expressions, occasionally does so with diminished intensity and fails to generate accurate lip motion for some words. A notable example is shown in the third column of the left case: when the speaker utters "um"; the expected closed-mouth motion is replaced by an open mouth in StyleTalk’s output. Besides, as shown in Fig. 4, StyleTalk occasionally generates videos where the face distorts suddenly.
Fig. 5 present comparisons with DiffTalk. DiffTalk struggles with lip synchronization and introduces jitteriness and artifacts in the mouth region.
@@
Figure 4: Face distortion observed in StyleTalk’s output.
Figure 5: Comparisons with DiffTalk.
Figure 6: The results generated using out-of-domain portraits.
Generalization Capabilities. Supplementary Video demonstrates our method’s capability to produce realistic talking head videos for out-of-domain portraits (paintings, toonified portraits, and sculptures, shown in Fig. 6), speech in various languages, noisy audio input, and songs. Sec. B.1 provides more analysis.
@@
Results of Speaking Style Prediction. Fig. 7 presents the results of speaking style predictions. The speakers with angry emotion are sampled from MEAD while those with fear emotion are sampled from RAVEDESS. The style predictor, utilizing emotional audio and neutral portraits from test videos, adeptly deduces personalized speaking styles that align with those observed in the original videos. It demonstrates the capacity to discern subtle expressions within the same emotion. For instance, for samples with angry emotion, the first-row speaker exhibits narrowed eyes, in contrast to the second-row speaker’s intense, glaring stare. For samples with angry emotion, the first-row speaker’s eyes and mouth are open, whereas the second-row speaker combines narrowed eyes with a contorted facial expression.
The ablation study and user study for the style predictor are presented in Appendix A.
4.3 Ablation Study
Figure 8: Ablation study results
@@ 4.3 Ablation Study
Table 2: The results of DreamTalk’s ablation study on MEAD. We omit CPBD scores since there are no significant differences between variants on CPBD.
To analyze the contributions of our designs, we conduct an ablation study with two variants: (1) remove the style-aware lip expert (w/o lip expert); (2) trained with unconditional lip expert (uncond lip expert). Our full model is denoted as Full.
Fig. 8 and Tab. 2 present our ablation study results. The variant w/o lip expert exhibits a decline in lip-sync accuracy on the emotional dataset MEAD, despite its competitive F-LMD score indicating expressive facial generation. Conversely, uncond lip expert secures a superior SyncNet confidence score at the expense of speaking style expressiveness. The Full model achieves a harmonious balance, ensuring both precise lip synchronization and vivid expressions, thanks to the style-aware lip expert directing the diffusion model’s expressive potential.
4.4 Style Code Visualization
Figure 9: t-SNE visualization of style codes from 15 speakers. Each color stands for style codes from an identical speaker.
Using t-distributed stochastic neighbor embedding (t-SNE) [75], we map style codes from the MEAD dataset’s 15 speakers into a 2D space. These speakers exhibit 22 distinct speaking styles, comprising seven emotions at three intensity levels, alongside a neutral style. For each style, style codes are extracted from 10 randomly chosen videos.
Fig. 9 reveals that style codes from identical speakers tend to cluster, despite some codes sharing identical emotion categories. This suggests that the variance in speaking styles attributable to individual speaker idiosyncrasies outweighs that induced by different emotions, thereby underpinning the rationale for using portrait information to infer speaking styles. We also observe that each speaker’s style code distribution exhibits both common patterns and individualized characteristics. We present these intriguing observations in Sec. B.2.
4.5 Style Manipulation
Figure 10: Effect of the input style controlled by adjusting the scale $\omega$ of the classifier-free guidance.
Adjusting the scale of Classifer-free Guidance. As elaborated in Sec. 3.3, the scale factor $\omega$ in the classifier-free guidance scheme modulates the effect of the input style. Adjusting $\omega$ either amplifies or attenuates the designated style, as shown in Fig. 10. When $\omega=0$, DreamTalk produces a talking head with a neutral expression. We observed that when the scale factor $\omega$ exceeds $2$, there is a noticeable decline in lip-sync accuracy.
Figure 11: The results of speaking style interpolation.
Style Code Interpolation. Leveraging the style space, we can modify speaking styles via style code manipulation. Fig. 11 illustrates that linear interpolation between style codes results in a seamless transition of generated speaking styles. This interpolation process allows for style intensity modulation and the generation of novel speaking styles.
4.6 User Study
[…]
MakeItTalk [105] 1.94 2.03 1.65
Wav2Lip [49] 2.29 1.45 1.14
PC-AVS [104] 2.26 1.81 1.86
AVCT [84] 2.31 2.21 1.72
GC-AVT [37] 2.35 1.20 1.88
EAMM [30] 1.81 1.40 1.78
StyleTalk [46] 2.35 2.29 2.08
SadTalker [100] 2.37 2.38 1.73
PD-FGC [78] 1.95 1.61 2.31
EAT [20] 2.24 1.65 2.29
DreamTalk 2.55 2.60 2.46
Ground Truth 3.03 2.89 1.83
Table 3: User study results.
We conduct a user study of 20 participants. We generate the test samples covering multiple speaking styles and speakers. For each method, the participant is required to score $10$ videos sampled from the test samples and is asked to give a rating (from 1 to 5, 5 is the best) on three aspects: (1) the Lip sync quality, (2) the realness of results, and (3) the style consistency between the generated videos and the style reference. As shown in Tab. 3, our method outperforms existing approaches across all aspects, particularly in style consistency, highlighting its superior capabilities.
5 Conclusion
In this work, we propose DreamTalk, a novel approach leveraging diffusion models for generating expressive talking heads. Our method aims to excel in diverse speaking styles while minimizing dependence on extra style references. We develop a denoising network for creating expressive, audio-driven facial motions and introduce a style-aware lip expert to optimize lip-sync without compromising style expressiveness. Additionally, we devise a style predictor that infers speaking styles directly from audio, eliminating the need for video references. The efficacy of DreamTalk is validated through extensive experiments.
Acknowledgements. This work is supported by Alibaba Group through Alibaba Research Intern Program. We would like to thank Xinya Ji, Borong Liang, Yan Pan, and Suzhen Wang for their generous help with the comparisons.
Supplementary Material
Appendix A Additional Evaluation for Speaking Style Prediction
A.1 Ablation Study
@@
To evaluate the impact of our design choices, we conduct an ablation study with three variants: (1) omitting speaker information and relying solely on audio for prediction (w/o speaker info); (2) during model training, the speaker info and audio are both obtained from the same video (w/o cross-ID training); (3) employing a regression model instead of a diffusion model for prediction (regression). Our full model is denoted as Full. When generating samples for evaluation, the facial images and audio we use are sourced from videos of the same individual expressing different emotions(e.g. the face image is from a happy video while the audio is from an angry one.). This generation approach better aligns with real-world applications.
@@
  • Style Accuracy (SA) We split test videos into different speaking styles and train a style classifier to classify which style a face motion sequence belongs to. Then, We classify the face motions generated using predicted style codes and report the accuracy. Specifically, we put the videos from the same speaker, emotion, and intensity into one style. We evaluate this metric on MEAD only since the number of RAVEDESS videos for each style is inadequate to train a style classifier. The ground truth testing set gets $92.5\%$ accuracy.
@@
Table 4: The ablation study results of the style predictor.
@@
The results are shown in Tab. 4 and Fig. 12. The w/o speaker info variant successfully predicts emotions from audio but occasionally fails to maintain consistency between the predicted speaking style and speaker identity, leading to poor identity preservation. This underscores the importance of speaker information in predicting speaking styles. Although in experiments, we observed that w/o cross-ID training achieves slightly better performance than Full when the input portrait and audio are from the same video, it underperformed, often failing to predict the correct emotion, when inputs were from different videos. This suggests that identity 3DMM parameters may convey some expression information, and without cross-ID training, the model might derive emotional cues from this leaked information rather than the audio. The regression variant struggles to generate accurate expressions for certain data, highlighting the superior distribution-learning capability of diffusion models in facilitating speaking style prediction.
A.2 User Study
In our user study, we evaluate the alignment between the original and predicted speaking styles. Directly assessing the alignment of speaking styles can be somewhat ambiguous, so we employ a comparative approach for evaluation. Specifically, we create a series of video triplets. Each triplet consisted of a test video from our dataset and two generated videos. The first video was generated using a style code predicted from an input portrait, sharing the same speaker identity as in the test video but displaying a neutral emotion, combined with the audio from the test video. The second video is generated using the style code extracted from videos with the same emotion but from a speaker different from the one in the test video. We recruit 20 participants. Each participant is then asked to evaluate 20 triplets and identify which of the generated videos most accurately reflected the speaking style of the test video. The videos generated using predicted style codes are preferred in $75.8\%$ of all ratings. This indicates that the style predictor is able to infer personalized speaking styles that are aligned with the audio.
A.3 Analyzing the Influence of Portraits
Appendix B Additional Results for Expressive Talking Head Generation
B.1 Analysis on Generalization Capabilities
Songs. As demonstrated in Supplementary Video, our method successfully generates reasonable results for songs, even those with accompaniment, despite this being significantly different from the training dataset’s data distribution. A noticeable decline in lip-sync accuracy is observed when the accompaniment volume is excessively high. We conduct a comparative analysis of lip-sync performance between songs with accompaniment and songs with removed accompaniment (using data from the SingFace Dataset [38]). It is found that the accompaniment adversely affects lip-sync, leading to mouth movements resembling mumbling. Addressing the negative impact of accompaniment on lip-sync accuracy presents an interesting avenue for future research.
Speech in Multiple Languages. Supplementary Video shows that our method generates satisfactory results with speech in French, Chinese, Spanish, German, Italian, Japanese, and Korean. The versatility of wav2vec features aids application across various languages. Additionally, the inclusion of multilingual talking head videos from Voxceleb2 enhances generalization.
Noisy Audio. Supplementary Video demonstrates that our method yields satisfactory outcomes when processing audio mixed with multiple noise types and intensities. We employ noise recordings from typical talking head application environments—meetings, offices, and cafeterias—sourced from the DEMAND dataset [73]. Using a publicly available tool [^1] https://github.com/Sato-Kunihiko/audio-SNR, we blend the audio with noise at SNRs of 20 dB, 10 dB, and 0 dB. Remarkably, our method maintains performance even at 0 dB SNR, where the noise is as loud as the speech and significantly impairs speech intelligibility and clarity.
B.2 More Results of Style Code Visualization

### v2

@@ III-C Training and Inference
instead of eq. 2. This approach enables controlling the effectiveness of the style reference $\bm{R}$ through adjustment of the scale factor $\omega$.
@@
We utilize PIRenderer [91] as the renderer and fine-tune it on emotional dataset to enable it to generate emotional talking faces.
Inference. Our method can specify speaking styles using either reference videos or solely through input audio and portrait. In the case of reference videos, style codes are derived using the style encoder in the denoising network. When relying solely on input audio and portrait, the style predictor takes these inputs and employs a denoising procedure to obtain the style code.
With the style code, the denoising network utilizes the sampling algorithm of DDPM [10] to produce face motions. It first samples a random motion $\bm{m}_{(T)}^{*}\sim{}\mathcal{N}(\bm{0},\bm{I})$ then computes denoised sequences $\{\bm{m}_{(t)}^{*}\},{t=T-1,\dots,0}$ by incrementally removing the noise from $\bm{m}_{(t)}^{*}$. Finally, the motion $\bm{m}_{(0)}^{*}$ is the generated face motion. The sampling process can be accelerated by leveraging DDIM [96]. The output face motions are then rendered into videos by the renderer.
IV Experiments
Figure 3: Qualitative comparisons. The red arrow indicates inaccurate lip motions.
IV-A Experimental Setup
@@
Baselines. We compare our method with previous methods: MakeitTalk [55], Wav2Lip [54], PC-AVS [59], AVCT [47], GC-AVT [3], EAMM [5], StyleTalk [2], DiffTalk [18], SadTalker [60], PD-FGC [1], and EAT [6]. DiffTalk’s released model cannot generate reasonable results until submission, so we perform qualitative comparisons using videos from its demo. For other methods, we generate the samples using released models or with authors’ help. When generating samples, we use the audio and the first image from the test video as inputs. We use a segment of the test video as the reference. Except when evaluating the style predictor, the style of DreamTalk is specified by the video.
Metrics. To evaluate video quality, we use SSIM [99] and the CPBD [100]. To evaluate lip-motion accuracy, we use the SyncNet confidence score ($\text{Sync}_{\text{conf}}$[101] and the Landmark Distance around mouth area (M-LMD) [53]. To evaluate the accuracy of generated expressions, we use the Landmark Distance on the full face (F-LMD) and a newly proposed metric Style Accuracy (SA). SA is the accuracy obtained from classifying samples using a speaking style classifier. When training the classifier, we divide the MEAD dataset into several groups with approximately consistent speaking styles and train the classifier to sort videos into the correct groups. Therefore, if a method generates accurate expressions, its samples will be classified into correct group and hence it will get higher SA. The details of SA metric are reported in Supp. Mat..
IV-B Main Results
@@
Quanitative Comparisons. As shown in table I, our method outperforms previous methods across most metrics. Wav2Lip’s SyncNet confidence score is higher than ours, even surpassing the ground truth. This is because Wav2Lip is trained using SyncNet as a discriminator. Notably, our method’s SyncNet confidence score closely aligns with the ground truth, and it achieves the best M-LMD scores, which indicates its capability for precise lip synchronization. Furthermore, our superior performance in the F-LMD and SA metrics demonstrates our method’s proficiency in generating facial expressions consistent with the reference speaking style.
Qualitative comparisons. fig. 3 shows the qualitative comparisons. The portraits, style references, and audio are all unseen during training.
Figure 4: Comparisons with StyleTalk on in-the-wild style reference. StyleTalk fails to generate accurate emotion.
Figure 5: Sudden face distortion (marked in red box) that frequently occurs in StyleTalk’s output. Better viewed in Supp. Video
TABLE II: Comparisons of identity preservation on MEAD. DreamTalk’s score is competitive to non-emotional methods and is the best in emotional methods.
Method MakeitTalk Wav2Lip PC-AVS AVCT SadTalker GCAVT EAMM StyleTalk PD-FGC EAT TH-PAD Ours Ground Truth
CSIM $\uparrow$ 0.77 0.89 0.17 0.79 0.78 0.20 0.36 0.64 0.35 0.61 0.30 0.71 0.80
StyleTalk, one of the most competitive baselines, fails to consistently generate high-quality results across diverse speaking styles. Firstly, StyleTalk frequently generates videos with sudden facial distortions (fig. 5), making it unsuitable for large-scale practical applications. We do not observe such phenomena in DreamTalk’s results. We speculate that the reason for Styletalk’s unstable generation is that StyleTalk uses overly complex loss functions and is based on GANs, models that suffer from unstable training. Secondly, StyleTalk fails to generate rich emotions, especially for in-the-wild style references. As shown in fig. 3, the output from StyleTalk exhibits discrepancies with the style reference: the left speaker’s eyes are not as narrowed, and the right speaker’s mouth is not opened as widely. The discrepancies are more pronounced when using in-the-wild style references. As shown in fig. 5, Styletalk fails to generate expressions consistent with those in the style reference, including raised eyebrows, glaring eyes, and widened mouths. Initially, we speculate that these issues are caused by the limited training data of StyleTalk, but we observe that when trained only using the same data as StyleTalk, DreamTalk can still generate expressions consistent with the style reference (the result is also shown in fig. 5). Therefore, we speculate that the reason is that GAN’s mode-collapse issue impairs the performance across diverse speaking styles. Thirdly, StyleTalk’s lip-sync is inferior. A notable example is shown in fig. 3: when the speaker utters "m"; the expected closed-mouth motion is replaced by an open mouth in StyleTalk’s output.
@@
It can be seen that MakeItTalk and AVCT struggle with accurate lip synchronization. While Wav2Lip and PC-AVS synchronize lips accurately, their outputs appear blurry. SadTalker, on the other hand, generally aligns lip movements with audio but occasionally displays unnatural jitters. EAT can only generate discrete emotions, lacking the finesse for nuanced expressions. For example, in the left case, the style reference shows the speaker narrowing his eyes, but EAT merely produces a generic disgusted look with wide-open glaring eyes. Additionally, as shown in the right case, EAT struggles to maintain a consistent face shape during speaker head movements. EAMM, GC-AVT, PD-FGC can produce fine-grained emotions. However, EAMM falls short in lip synchronization, GC-AVT and PD-FGC struggle with preserving speaker identity, and all three have issues rendering a plausible background. As shown in fig. 7, DiffTalk struggles with lip synchronization and produces jitteriness and artifacts in the mouth area.
@@
Figure 6: Comparisons with DiffTalk. DiffTalk fails to achieve lip-sync and produces jitteriness.
Figure 7: Emotional results generated from songs in multiple languages (English, Chinese, Japanese).
@@
Evaluation of the Identity Preservation. To evaluate the ability to preserve identity, we utilize a widely used metric CSIM. When computing the CSIM score, we utilize an off-the-shelf face recognition network ArcFace [102] to extract the deep identity features from each generated frame and then calculate the cosine similarities between the features of the input portrait and the generated frames. We find that when the ID in the style reference and portrait differ significantly, ID preservation worsens. So, we compute scores for StyleTalk and DreamTalk when the style reference comes from a randomly selected ID different from that in portrait. table II shows the result on MEAD dataset. Wav2Lip attains the highest score since it merely changes the mouth region and leaves other parts intact. Non-emotional methods achieve better scores than emotional ones because changing emotions may change the identity perceived by humans and ArcFace. DreamTalk’s score is competitive to non-emotional methods and is the best in emotional methods.
Generalization Capabilities. Leveraging the strong power of diffusion models, DreamTalk shows strong generalization capabilities. As shown in fig. 7 and Supp. Video, DreamTalk can even generate reasonable results for songs in multiple languages, even though our training dataset includes only a few multilingual audio and no song audio. Supp. Video shows that DreamTalk also generalizes well to speech in various languages and noisy audio.
Results of Speaking Style Prediction. fig. 8 presents the results of speaking style predictions. The style predictor, utilizing emotional audio and neutral portraits, adeptly deduces personalized speaking styles aligned with those in the original videos. It can discern subtle expressions within the same emotion. For instance, for samples with angry emotion, the first-row speaker exhibits narrowed eyes, in contrast to the second-row speaker’s intense, glaring stare. For samples with fear emotion, the first-row speaker’s eyes and mouth are open, whereas the second-row speaker combines narrowed eyes with a contorted facial expression.
Figure 10: Comparisons with TH-PAD.
Figure 11: Ablation study results of emotional talking head generation.
We compare the style predictor with TH-PAD [19], a method that also uses audio to predict expressions. We obtain TH-PAD samples with the authors’ help. For comparisons, we use audio to predict speaking styles and use predicted styles to generate samples. We observe that TH-PAD fails to generate emotions conveyed in audio. TH-PAD only generates neutral expressions aligned with the audio rhythm. fig. 11 shows that TH-PAD fails to generate the sad emotion reflected in the audio. We also conduct a quantitative comparison on MEAD. As shown in table III, TH-PAD’s SA and LMD scores, which measure emotion alignment, are inferior.
TABLE III: Comparisons with TH-PAD on MEAD. DreamTalk(A) uses styles predicted from audio.
[…]
       TH-PAD        6.38/5.81        5.3
       DreamTalk(A)        2.24/3.43        78.6
       Ground Truth        0 / 0        92.5
TABLE IV: The results of DreamTalk’s ablation study on MEAD. CPBD is omitted due to no significant differences.
@@
IV-C Ablation Study
Emotional Talking Head Generation. To analyze the contributions of our designs, we conduct an ablation study with two variants: (1) remove the style-aware lip expert (w/o lip expert); (2) trained with unconditional lip expert (uncond lip expert). Our full model is denoted as Full.
fig. 11 and table IV present our ablation study results. The variant w/o lip expert exhibits a decline in lip-sync accuracy on the emotional dataset MEAD, despite its competitive F-LMD score indicating expressive facial generation. Conversely, uncond lip expert secures a superior SyncNet confidence score at the expense of speaking style expressiveness. The Full model achieves a harmonious balance, ensuring both precise lip synchronization and vivid expressions, thanks to the style-aware lip expert directing the diffusion model’s expressive potential.
@@ IV-C Ablation Study
Speaking Style Prediction. To evaluate the impact of our design choices, we conduct an ablation study with three variants: (1) omitting speaker information and relying solely on audio for prediction (w/o speaker info); (2) during model training, the speaker info and audio are both obtained from the same video (w/o cross-ID training); (3) employing a regression model instead of a diffusion model for prediction (regression). Our full model is denoted as Full. When generating samples for evaluation, the facial images and audio we use are sourced from videos of the same individual expressing different emotions(e. g. the face image is from a happy video while the audio is from an angry one.). This generation approach better aligns with real-world applications.
@@
  • Style Accuracy (SA) We use the SA metric mentioned in the section IV-A. SA is evaluated on 3DMM face motions. The ground truth testing set gets $92.5\%$ accuracy.
@@
TABLE V: The ablation study results of the style predictor.
@@
The results are shown in table V and fig. 12. The w/o speaker info variant successfully predicts emotions from audio but occasionally fails to maintain consistency between the predicted speaking style and speaker identity, leading to poor identity preservation. This underscores the importance of speaker information in predicting speaking styles. Although in experiments, we observed that w/o cross-ID training achieves slightly better performance than Full when the input portrait and audio are from the same video, it underperformed, often failing to predict the correct emotion, when inputs were from different videos. This suggests that identity 3DMM parameters may convey some expression information, and without cross-ID training, the model might derive emotional cues from this leaked information rather than the audio. The regression variant struggles to generate accurate expressions for certain data, highlighting the superior distribution-learning capability of diffusion models in facilitating speaking style prediction.
IV-D Style Code Visualization